NvidiaIntely (nvidiaintely):

In the reaction, CH4 + 2O2 arrow CO2 + 2H2O, 890 kilojoules of heat are released. Using Einstein's equation, calculate the mass amount to which this energy equates. (1 J = 1 kg m2 s-2) and the speed of light = 3.0 x 108 m/sec 9.9 x 1012 kg 9.9 x 10-12 kg 9.9 x 10-14 kg 1.1 x 10-6 kg

NvidiaIntely (nvidiaintely):

@Kevin @Will.H

NvidiaIntely (nvidiaintely):

Oh already here lol

OpenStudy (will.h):

9.9 x 10-12 kg

OpenStudy (jiteshmeghwal9):

E=mc^2 E=890×10^3 joules C=3×10^8 m/s

OpenStudy (jiteshmeghwal9):

M=?

OpenStudy (will.h):

m = e/c^2